Cross-price elasticities

Rice & beans (-0.35)

Rice & wheat (0.6)

Rice & chicken (-0.1)

Rice & milk (-0.05)

Rice & other goods 0

Income elasticity of demand for rice (0.4)

Questions:

D) How much would the price of rice have to decrease in order to increase rice consumption by 5%?

E) What would happen to bean consumption as a result of a 10 percent decrease in the price of rice? (Make sure to mention the direction and magnitude of the impact.)

F) What would happen to chicken consumption as a result of a 10 percent increase in the price of rice? (Make sure to mention the direction and magnitude of the impact.)
Answer #1

D)

Own-price elasticity of demand for rice is not given. Assume it to be -0.3

Thus, when rice consumption increase by 5%

The % decrease in the price of the rice is 5% of (0.3) = 1.5%

`E) The cross price elasticity of Rice & beans is -0.35

When % decrease in the price of the rice is 10%

The % increase in the consumption of beans = 10*0.35 = 3.5%

Thus, the magnitude of increase in beans consumption = 3.5%

The direction of change in beans consumption is increase in beans consumption,following an decrease in price of the rice

F)

The cross price elasticity of Rice & chickens is -0.1

When % increase in the price of the rice is 10%

The % decrease in the consumption of chickens = 10*0.1 = 1%

Thus, the magnitude of decrease in chickens consumption = 1%

The direction of change in chickens consumption is decrease in chicken consumption following an increase in price of the rice